# Weight Space Representation Learning on Diverse NeRF Architectures

### v2

@@ 4 Experiments
Datasets. Our experimental evaluation is based on three families of NeRF architectures: MLP-based (i.e. a vanilla MLP), tri-planar (i.e. a tri-plane [8] followed by an MLP), and hash-based (i.e. a multi-resolution hash table [27] followed by an MLP). These families are exemplified by the following datasets: (i) one consisting of Zama Ramirez et al. [38]’s MLP-based NeRFs only, which will be referred to as MLP; (ii) one consisting of Cardace et al. [7]’s tri-planar NeRFs only, which will be referred to as TRI; (iii) one, created by us, consisting of hash-based NeRFs only, which will be referred to as HASH; (iv) the union of MLP, TRI, and HASH, i.e. a dataset where each object appears three times: once as an MLP-based NeRF, then as a tri-planar NeRF, and finally as a hash-based NeRF; we will refer to this dataset as ALL. To further assess the ability of our framework to perform tasks on arbitrary NeRF architectures, we create additional test sets of NeRFs featuring architectures belonging to the same families of those seen at training time but with different hyperparameters, which we will refer to as unseen architectures. The specifics of the architectures of all our datasets are reported in Table 1. Overall, we consider a total of 13 architectures across the three families. Following the protocol first introduced by Zama Ramirez et al. [38], all the above NeRFs are trained on ShapenetRender [37], a dataset providing RGB images of synthetic 3D objects together with their class label.
[…]
Single vs multi-architecture. Our method is the first neural processing framework able to handle multiple NeRF architectures. Thus, the natural setting to test it is when it is trained on ALL; we will refer to this scenario as the multi-architecture setting. Yet, our method can also be used when the input NeRFs all share the same architecture, e.g. when our framework has been trained on MLP, TRI, or HASH only, by dropping the contrastive loss term. Therefore, to test its generality, we also perform experiments in such a scenario, which we will refer to as the single-architecture setting. In this setting, our approach can be compared against previous methods [38, 7] that can only handle specific architectures.
@@ 4.1 Latent Space Analysis
Figure 3: t-SNE plots. 2D projections of the latent space created by our framework when trained on a dataset of NeRFs of ShapenetRender objects [37], where each object is represented by three NeRFs parameterized by different architectures: MLPs, tri-planes, and multi-resolution hash tables.
@@
To study the impact of different learning objectives on the organization of the resulting NeRF latent space, we apply t-SNE dimensionality reduction [33] to the embeddings computed by our framework on NeRFs belonging to the test set of the ALL dataset. Fig. 3 shows the resulting bi-dimensional plots. Some interesting patterns can be noted. When trained to minimize $\mathcal{L}_{\text{R}}$ alone (Fig. 3, left), the encoder creates an embedding space that is clustered by class, even though class labels were not used at training time. This behavior is a byproduct of the loss, which enforces NeRFs encoding similar shapes and colors to lie nearby in the latent space. Another outcome one may expect is that the same object represented by different NeRF architectures is projected into nearby embeddings; however, this turns out not to be the case. Instead, three distinct clusters emerge for each class, each corresponding to a NeRF architecture, and these clusters are further away from each other than clusters sharing the same underlying architecture. This outcome shows that $\mathcal{L}_{\text{R}}$ alone does not directly encourage the model to align NeRF embeddings regardless of the input architecture. Conversely, when trained to minimize $\mathcal{L}_{\text{C}}$ alone (Fig. 3, right), the encoder greatly reduces the distance between different architectures representing the same class, but also provides a significantly less distinguishable clusterization by class compared to $\mathcal{L}_{\text{R}}$ case (especially for some classes, like chairs, sofas, and tables). Finally, $\mathcal{L}_{\text{R}+\text{C}}$ (Fig. 3, middle) strikes a balance between these latent space properties. Compared to $\mathcal{L}_{\text{R}}$, one macro-cluster per class is present, although hash tables have a tendency to form a separate sub-cluster (which is especially evident for airplanes and watercrafts). Compared instead to $\mathcal{L}_{\text{C}}$, classes are more separated (although not as much as with $\mathcal{L}_{\text{R}}$), but the sub-clusterization of hash tables is more apparent. Thus, we expect $\mathcal{L}_{\text{R}}$ to be the best choice for tasks where the separation between classes is highly relevant; $\mathcal{L}_{\text{R}+\text{C}}$ and $\mathcal{L}_{\text{C}}$, on the other hand, are likely to be the most effective for tasks where strong invariance to the NeRF architecture is required, with $\mathcal{L}_{\text{R}+\text{C}}$ being preferable, as it preserves more separation between classes. For these reasons, the remainder of this section shows experimental results obtained with $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$; results for $\mathcal{L}_{\text{C}}$ are reported in Appendix C.
@@
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) ALL 93.6 94.0 92.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 90.7 90.6 90.0
$\mathcal{L}_{\text{R}}$ (ours) MLP 93.8 25.3 19.3
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 91.5 58.6 56.6
$\mathcal{L}_{\text{R}}$ (ours) TRI 11.4 93.8 9.3
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 77.8 91.0 66.5
$\mathcal{L}_{\text{R}}$ (ours) HASH 13.8 35.7 92.7
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 54.1 35.5 90.6
@@ 4.2 NeRF Classification
Once our framework has been trained, NeRF classification is performed by learning a downstream classifier $\mathcal{C}$ to predict the labels of the embeddings produced by the encoder; an overview of this procedure is shown in Fig. 1. We report results for two cases: classification of NeRF architectures on which our framework was trained and classification of unseen architectures. In both cases, since the multi-architecture setting requires evaluating methods trained on ALL, neither one of the previous works, i.e. nf2vec [38] and Cardace et al. [7], can be applied to this scenario, as they can only be trained on MLP and TRI, respectively.
Training architectures. Table 3 shows NeRF classification results in the multi-architecture setting on architectures used to train our framework. When $\mathcal{C}$ has been trained on ALL, $\mathcal{L}_{\text{R}}$ performs the best: this result is consistent with the better separation between clusters corresponding to different classes provided by $\mathcal{L}_{\text{R}}$, as shown in Fig. 3 (left) and discussed in Section 4.1. This trend can also be observed, as expected, when $\mathcal{C}$ is trained on either MLP, TRI, or HASH and tested on that same dataset, as this scenario replicates the previous one with smaller datasets. On the other hand, the introduction of a contrastive objective in $\mathcal{L}_{\text{R}+\text{C}}$ is key to performance whenever $\mathcal{C}$ is trained on a single-architecture dataset (e.g. MLP) and tested on a dataset with a different architecture (e.g. TRI or HASH), as clusters of NeRFs belonging to the same class but parameterized by different architectures are closer in the embedding space than with $\mathcal{L}_{\text{R}}$ alone. Overall, our method consistently achieves remarkable accuracies when tested on architectures included in the training set of $\mathcal{C}$ (i.e. red cells in Table 3). In the single-architecture setting, instead, previous methods can be trained and evaluated on the datasets corresponding to the NeRF architecture they were designed to process, whereas $\mathcal{L}_{\text{R}+\text{C}}$ cannot be applied, as there is a single architecture and, therefore, no positive pairs to compute the contrastive loss. As shown in Table 3, when both the encoder and the classifier are trained and tested on either MLP or TRI, our method outperforms both nf2vec and Cardace et al., while also achieving high accuracy on HASH. Hence, beyond its original aim, our approach can also provide a competitive alternative to single-architecture frameworks.
@@
Training Set
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) ALL 91.3 87.4 93.2 88.3 24.8 69.4 91.9 91.2 88.3 24.5
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 85.9 83.8 87.0 84.1 72.1 30.8 89.2 87.4 86.8 27.8
$\mathcal{L}_{\text{R}}$ (ours) MLP 91.3 86.1 22.6 23.2 7.9 21.0 19.7 20.6 21.8 7.6
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 86.6 81.3 63.3 43.4 59.2 13.1 51.5 50.8 55.0 24.8
$\mathcal{L}_{\text{R}}$ (ours) TRI 11.9 10.6 92.1 84.8 32.1 43.6 9.2 8.8 15.0 5.2
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 60.7 58.0 86.9 83.2 63.9 30.2 59.8 61.5 66.3 28.5
$\mathcal{L}_{\text{R}}$ (ours) HASH 10.7 7.4 33.9 36.3 19.1 23.0 91.6 91.4 87.8 29.2
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 47.0 40.4 40.6 34.0 43.4 19.3 89.5 87.9 86.4 25.7
@@
Method Encoder MLP-2L MLP-32H TRI-2L TRI-32H TRI-16W TRI-8C HASH-2L HASH-32H HASH-3N HASH-11T
Training Set
nf2vec [38] MLP 63.7
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) 91.8 83.7 10.9 6.1 7.0 4.9 5.7 5.6 4.0 4.9
nf2vec [38] TRI
Cardace et al. [7] 93.5 92.0 68.6
$\mathcal{L}_{\text{R}}$ (ours) 10.1 10.0 92.6 82.5 22.9 72.8 17.4 13.6 10.8 12.1
nf2vec [38] HASH
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) 8.3 8.9 15.4 21.3 21.3 21.3 90.6 91.0 88.2 28.8
Unseen architectures. Multi-architecture classification results on unseen architectures are shown in Table 4. Remarkably, a trend coherent with that of Table 3 can be noticed, namely that $\mathcal{L}_{\text{R}}$ tends to perform best when the classifier $\mathcal{C}$ is trained either on ALL or on an architecture belonging to the family of the unseen one presented at test time (i.e. orange cells in Table 4), whereas $\mathcal{L}_{\text{R}+\text{C}}$ almost always prevails when training and test families differ, usually by a significant margin. Notable exceptions are TRI-16W, where $\mathcal{L}_{\text{R}+\text{C}}$ consistently yields the best results, and TRI-8C, where $\mathcal{L}_{\text{R}}$ outperforms (although less dramatically) $\mathcal{L}_{\text{R}+\text{C}}$, even when changing architecture family. Some other interesting patterns can be noted: in MLP and tri-plane families, reducing the number of hidden layers of an MLP has less impact on accuracy than reducing the hidden dimension, whereas it does not seem to have a significant effect on hash table families; furthermore, changing hash table size is the most disruptive change, even when encoder and classifier have seen hash-based NeRFs at training time. These patterns can be the basis of future investigations into this topic. Overall, it is worth highlighting that, in the most relevant scenario, i.e. when our framework has been trained on ALL, our method is able to reach remarkable accuracies on most of the unseen variations of known architecture families it is tested on. In the single-architecture setting, shown in Table 5, $\mathcal{L}_{\text{R}}$ performs much better than nf2vec in the only case in which the latter can be tested. Conversely, $\mathcal{L}_{\text{R}}$ performs worse than Cardace et al. [7] in TRI-2L and TRI-32H, i.e. TRI variations that change the MLP architecture. This is not surprising since, when performing classification, Cardace et al. processes the tri-plane alone and discards the MLP, which makes their approach invariant to changes in the MLP architecture. Yet, since their method is tied to the tri-plane quantization, it cannot process TRI-16W and is less robust than ours to the reduction in the number of channels. Finally, our method is the only one that can be tested on the hash table family, where it exhibits robust performance when trained on HASH, with the exception of HASH-11T, which already emerged as particularly challenging in the results of Table 4.
Table 6: Instance-level NeRF retrieval. The encoder is trained on ALL. The reported query/gallery combinations belong to their corresponding test sets.
@@
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) 1 0.0 1.8 0.4 4.5 1.7 1.1 0.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 5.3 30.6 14.8 33.3 13.2 13.5 9.5
$\mathcal{L}_{\text{R}}$ (ours) 5 0.0 5.3 1.8 13.3 6.1 3.8 1.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 17.7 59.4 38.0 61.2 33.4 32.1 25.1
$\mathcal{L}_{\text{R}}$ (ours) 10 0.0 8.2 3.1 19.2 9.2 6.1 2.7
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 27.2 72.7 51.3 72.6 45.0 42.8 36.3
Figure 4: NeRF retrieval qualitative results ($\mathcal{L}_{\text{R}+\text{C}}$). Top: query and gallery from MLP, TRI, or HASH. Bottom: query from ObjaNeRF [3], gallery from MLP, TRI, or HASH.

### v3

@@ 4 Experiments
Datasets. Our experimental evaluation is based on three families of NeRF architectures: MLP-based (i.e. a vanilla MLP), tri-planar (i.e. a tri-plane (Chan et al., 2022) followed by an MLP), and hash-based (i.e. a multi-resolution hash table (Müller et al., 2022) followed by an MLP). These families are exemplified by the following datasets: (i) one consisting of Zama Ramirez et al. (2024)’s MLP-based NeRFs only, which will be referred to as MLP; (ii) one consisting of Cardace et al. (2024)’s tri-planar NeRFs only, which will be referred to as TRI; (iii) one, created by us, consisting of hash-based NeRFs only, which will be referred to as HASH; (iv) the union of MLP, TRI, and HASH, i.e. a dataset where each object appears three times: once as an MLP-based NeRF, then as a tri-planar NeRF, and finally as a hash-based NeRF; we will refer to this dataset as ALL. To further assess the ability of our framework to perform tasks on arbitrary NeRF architectures, we create additional test sets of NeRFs featuring architectures belonging to the same families of those seen at training time but with different hyperparameters, which we will refer to as unseen architectures. The specifics of the architectures of all our datasets are reported in Tab.˜1. Overall, we consider a total of 13 architectures across the three families. Following the protocol first introduced by Zama Ramirez et al. (2024), all the aforementioned NeRFs are trained on ShapenetRender (Xu et al., 2019), a dataset providing RGB images of synthetic 3D objects together with their class label.
Models. To assess the impact of different losses on training, we introduce a distinction between three versions of our framework, depending on the learning objective: (i) $\mathcal{L}_{\text{R}}$, where the framework has been trained with the rendering loss of Eq.˜2 alone; (ii) $\mathcal{L}_{\text{R}+\text{C}}$, where the framework has been trained with a combination of rendering and contrastive losses as in Eq.˜4, with $\lambda=$2\text{\times}{10}^{-2}$\$ (as it leads to similar magnitudes in the two terms); (iii) $\mathcal{L}_{\text{C}}$, where the framework has been trained with the contrastive loss of Eq.˜3 alone.
Single vs multi-architecture. Our method is the first neural processing framework able to handle multiple NeRF architectures. Thus, the natural setting to test it is when it is trained on ALL; we will refer to this scenario as the multi-architecture setting. Yet, our method can also be used when the input NeRFs all share the same architecture, e.g. when our framework has been trained on MLP, TRI, or HASH only, by dropping the contrastive loss term. Therefore, to test its generality, we also perform experiments in such a scenario, which we will refer to as the single-architecture setting. In this setting, our approach can be compared to previous methods (Zama Ramirez et al., 2024; Cardace et al., 2024) that are limited to handling specific architectures.
@@ 4.1 Latent Space Analysis
Figure 3: t-SNE plots. 2D projections of the latent space created by our framework when trained on a dataset of NeRFs of ShapenetRender objects (Xu et al., 2019), where each object is represented by three NeRFs parameterized by different architectures: MLPs, tri-planes, and multi-resolution hash tables.
@@
To study the impact of different learning objectives on the organization of the resulting NeRF latent space, we apply t-SNE dimensionality reduction (Van der Maaten & Hinton, 2008) to the embeddings computed by our framework on NeRFs belonging to the test set of the ALL dataset. Fig.˜3 shows the resulting bi-dimensional plots. Some interesting patterns can be noted. When trained to minimize $\mathcal{L}_{\text{R}}$ alone (Fig.˜3, left), the encoder creates an embedding space that is clustered by class, even though class labels were not used at training time. This behavior is a byproduct of the loss, which enforces NeRFs encoding similar shapes and colors to lie nearby in the latent space. Another outcome one may expect is that the same object represented by different NeRF architectures is projected into nearby embeddings; however, this turns out not to be the case. Instead, three distinct clusters emerge for each class, each corresponding to a NeRF architecture, and these clusters are further away from each other than clusters sharing the same underlying architecture. This outcome shows that $\mathcal{L}_{\text{R}}$ alone does not directly encourage the model to align NeRF embeddings regardless of the input architecture. Conversely, when trained to minimize $\mathcal{L}_{\text{C}}$ alone (Fig.˜3, right), the encoder significantly reduces the distance between different architectures representing the same class, but also provides a considerably less distinguishable clusterization by class compared to $\mathcal{L}_{\text{R}}$ case (especially for some classes, like chairs, sofas, and tables). Finally, $\mathcal{L}_{\text{R}+\text{C}}$ (Fig.˜3, middle) strikes a balance between these latent space properties. Compared to $\mathcal{L}_{\text{R}}$, one macro-cluster per class is present, although hash tables tend to form a separate sub-cluster (which is especially evident for airplanes and watercrafts). Compared to $\mathcal{L}_{\text{C}}$, classes are more separated (although not as much as with $\mathcal{L}_{\text{R}}$), but the sub-clusterization of hash tables is more apparent. Thus, we expect $\mathcal{L}_{\text{R}}$ to be the best choice for tasks where the separation between classes is highly relevant; $\mathcal{L}_{\text{R}+\text{C}}$ and $\mathcal{L}_{\text{C}}$, on the other hand, are likely to be the most effective for tasks where strong invariance to the NeRF architecture is required, with $\mathcal{L}_{\text{R}+\text{C}}$ being preferable, as it preserves more separation between classes. For these reasons, the remainder of this section shows experimental results obtained with $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$; results for $\mathcal{L}_{\text{C}}$ are reported in App.˜C.
@@
nf2vec
Cardace et al.
[…]
@@ 4.2 NeRF Classification
Once our framework has been trained, NeRF classification is performed by learning a downstream classifier $\mathcal{C}$ to predict the labels of the embeddings produced by the encoder; an overview of this procedure is shown in Fig.˜1. We report results for two cases: classification of NeRF architectures on which our framework was trained and classification of unseen architectures. In both cases, since the multi-architecture setting requires evaluating methods trained on ALL, neither one of the previous works, i.e. nf2vec (Zama Ramirez et al., 2024) and Cardace et al. (2024), can be applied to this scenario, as they can only be trained on MLP and TRI, respectively.
Training architectures. Tab.˜3 shows NeRF classification results in the multi-architecture setting on architectures used to train our framework. When $\mathcal{C}$ has been trained on ALL, $\mathcal{L}_{\text{R}}$ performs the best: this result is consistent with the better separation between clusters corresponding to different classes provided by $\mathcal{L}_{\text{R}}$, as shown in Fig.˜3 (left) and discussed in Sec.˜4.1. This trend can also be observed, as expected, when $\mathcal{C}$ is trained on either MLP, TRI, or HASH and then tested on the same dataset, replicating the previous scenario with smaller datasets. On the other hand, the introduction of a contrastive objective in $\mathcal{L}_{\text{R}+\text{C}}$ is key to performance whenever $\mathcal{C}$ is trained on a single-architecture dataset (e.g. MLP) and tested on a dataset with a different architecture (e.g. TRI or HASH), as clusters of NeRFs belonging to the same class but parameterized by different architectures are closer in the embedding space than with $\mathcal{L}_{\text{R}}$ alone. Overall, our method consistently achieves remarkable accuracies when tested on architectures included in the training set of $\mathcal{C}$ (i.e. red cells in Tab.˜3). In the single-architecture setting, instead, previous methods can be trained and evaluated on the datasets corresponding to the NeRF architecture they were designed to process, whereas $\mathcal{L}_{\text{R}+\text{C}}$ cannot be applied, as there is a single architecture and, therefore, no positive pairs to compute the contrastive loss. As shown in Tab.˜3, when both the encoder and the classifier are trained and tested on either MLP or TRI, our method outperforms both nf2vec and Cardace et al., while also achieving high accuracy on HASH. Hence, beyond its original aim, our approach can also provide a competitive alternative to single-architecture frameworks.
@@
Method Classifier Training Set MLP-2L MLP-32H TRI-2L TRI-32H TRI-16W TRI-8C HASH-2L HASH-32H HASH-3N HASH-11T
nf2vec
Cardace et al.
[…]
@@
Method Encoder Training Set MLP-2L MLP-32H TRI-2L TRI-32H TRI-16W TRI-8C HASH-2L HASH-32H HASH-3N HASH-11T
[…]
Unseen architectures. Multi-architecture classification results on unseen architectures are shown in Tab.˜4. Remarkably, a trend coherent with that of Tab.˜3 can be noticed, namely that $\mathcal{L}_{\text{R}}$ tends to perform best when the classifier $\mathcal{C}$ is trained either on ALL or on an architecture belonging to the family of the unseen one presented at test time (i.e. orange cells in Tab.˜4), whereas $\mathcal{L}_{\text{R}+\text{C}}$ almost always prevails when training and test families differ, usually by a significant margin. Notable exceptions are TRI-16W, where $\mathcal{L}_{\text{R}+\text{C}}$ consistently yields the best results, and TRI-8C, where $\mathcal{L}_{\text{R}}$ outperforms $\mathcal{L}_{\text{R}+\text{C}}$ (although less dramatically) even when changing the architecture family. Some other interesting patterns can be noted: in MLP and tri-plane families, reducing the number of hidden layers of an MLP has less impact on accuracy than reducing the hidden dimension, whereas it does not seem to have a significant effect on hash table families; furthermore, changing hash table size is the most disruptive change, even when encoder and classifier have seen hash-based NeRFs at training time. These patterns can serve as the basis for future investigations into this topic. Overall, it is worth highlighting that, in the most relevant scenario, i.e. when our framework has been trained on ALL, our method is able to reach remarkable accuracies on most of the unseen variations of known architecture families it is tested on. In the single-architecture setting, shown in Tab.˜5, $\mathcal{L}_{\text{R}}$ performs much better than nf2vec in the only case in which the latter can be tested. Conversely, $\mathcal{L}_{\text{R}}$ performs worse than Cardace et al. (2024) in TRI-2L and TRI-32H, i.e. TRI variations that change the MLP architecture. This is not surprising since, when performing classification, Cardace et al. processes the tri-plane alone and discards the MLP, which makes their approach invariant to changes in the MLP architecture. Yet, since their method is tied to the tri-plane quantization, it cannot process TRI-16W and is less robust than ours to the reduction in the number of channels. Finally, our method is the only one that can be tested on the hash table family, where it exhibits robust performance when trained on HASH, except for HASH-11T, which already emerged as particularly challenging in the results of Tab.˜4.
Table 6: NeRF retrieval (ShapenetRender). The encoder is trained on ALL. Query/gallery combinations belong to their respective test sets.
@@
Method $k$ MLP/TRI MLP/HASH TRI/MLP TRI/HASH HASH/MLP HASH/TRI
nf2vec
Cardace et al.
Random 1 0.03 0.03 0.03 0.03 0.03 0.03
$\mathcal{L}_{\text{R}}$ (ours) 1.80 0.43 4.48 1.67 1.06 0.38
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 30.62 14.77 33.27 13.17 13.45 9.46
Random 5 0.13 0.13 0.13 0.13 0.13 0.13
$\mathcal{L}_{\text{R}}$ (ours) 5.28 1.82 13.30 6.12 3.84 1.37
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 59.37 37.98 61.24 33.40 32.09 25.13
Random 10 0.25 0.25 0.25 0.25 0.25 0.25
$\mathcal{L}_{\text{R}}$ (ours) 8.19 3.08 19.24 9.23 6.09 2.68
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 72.67 51.25 72.62 45.01 42.83 36.31
Table 7: NeRF retrieval (Objaverse generalization). The encoder is trained on ALL. Query/gallery combinations belong to their respective test sets.
Recall@$k$ (%) $\uparrow$
Method $k$ $\texttt{MLP}^{\texttt{OB}}$$\texttt{TRI}^{\texttt{OB}}$ $\texttt{MLP}^{\texttt{OB}}$$\texttt{HASH}^{\texttt{OB}}$ $\texttt{TRI}^{\texttt{OB}}$$\texttt{MLP}^{\texttt{OB}}$ $\texttt{TRI}^{\texttt{OB}}$$\texttt{HASH}^{\texttt{OB}}$ $\texttt{HASH}^{\texttt{OB}}$$\texttt{MLP}^{\texttt{OB}}$ $\texttt{HASH}^{\texttt{OB}}$$\texttt{TRI}^{\texttt{OB}}$
nf2vec
Cardace et al.
Random 1 0.05 0.05 0.05 0.05 0.05 0.05
$\mathcal{L}_{\text{R}}$ (ours) 2.60 0.48 2.71 0.48 0.74 0.32
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 9.82 6.10 8.33 1.75 6.74 2.92
Random 5 0.27 0.27 0.27 0.27 0.27 0.27
$\mathcal{L}_{\text{R}}$ (ours) 6.37 1.11 7.27 1.91 2.92 0.69
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 23.67 17.94 20.70 5.73 19.59 8.49
Random 10 0.53 0.53 0.53 0.53 0.53 0.53
$\mathcal{L}_{\text{R}}$ (ours) 9.66 2.39 10.99 3.24 4.88 1.80
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 33.65 26.38 27.55 9.50 28.13 12.74
Figure 4: Across-datasets NeRF retrieval ($\mathcal{L}_{\text{R}+\text{C}}$). Query from the test set of $\texttt{MLP}^{\texttt{OB}}$ (Amaduzzi et al., 2025), gallery from the test set of MLP, TRI, or HASH.